OpenStudy (anonymous):

find a,b,c for which the matrix (top: a, 1/sqr2, -1/sqr2; middle: b,1/sqr6, 1/sqr6; bottom: c, 1/sqr3, 1/sqr3 ) is orthogonal. Are the values of a, b,c unique? Explain. Help, please

OpenStudy (anonymous):

\[\left[\begin{matrix}a & \frac{ 1 }{ \sqrt{2} }& \frac{ -1 }{ \sqrt{2} } \\ b & \frac{ 1 }{ \sqrt{6} }& \frac{ 1 }{ \sqrt{6} }\\c & \frac{ 1 }{ \sqrt{3} }& \frac{ 1 }{ \sqrt{3} }\end{matrix}\right]\]

OpenStudy (anonymous):

the columns should be orthogonal unit vectors.. so, the last two columns: \(-{1\over2}+{1\over6}+{1\over3}=0\) and \({1\over2}+{1\over6}+{1\over3}=1\) so, the last two columns are orthogonal unit vectors.

OpenStudy (anonymous):

the first column has to be (1) a unit vector so, \(\qquad a^2+b^2+c^2=1\) (2) orthogonal to column 2 so, \(\qquad {a\over\sqrt{2}}+{b\over\sqrt{6}}+{c\over\sqrt{3}}=0\) (3) orthogonal to column 3 so, \(\qquad {-a\over\sqrt{2}}+{b\over\sqrt{6}}+{c\over\sqrt{3}}=0\)

OpenStudy (anonymous):

you mean orthonormal?

OpenStudy (anonymous):

eq (2) - eq(3) gives: \(\large2{a\over\sqrt{2}}=0\implies a=0\)

OpenStudy (anonymous):

then we have \[b^2=1-c^2\\ {\rm and}\qquad{b\over\sqrt{6}}=-{c\over\sqrt{3}}\implies b^2=2c^2\\ \therefore 3c^2=1\implies \boxed{\large\color{red} {c=\pm{1\over\sqrt{3}}}}\]

OpenStudy (anonymous):

and \[\boxed{\large\color{red} {b=\pm{\sqrt{2\over3}}}}\]

OpenStudy (anonymous):

so, there is no unique asnwer but "4" possible solutions

OpenStudy (anonymous):

\[(a,b,c) = (0,\sqrt{2/3},1/\sqrt{3})\\ (a,b,c) = (0,\sqrt{2/3},-1/\sqrt{3})\\ (a,b,c) = (0,-\sqrt{2/3},1/\sqrt{3})\\ (a,b,c) = (0,-\sqrt{2/3},-1/\sqrt{3}) \]

OpenStudy (anonymous):

now, which of these gives the orthogonal componetns?

OpenStudy (anonymous):

@Hoa you follow?

OpenStudy (anonymous):

we notice that when "b" is positive, "c" HAS to be negative and vice-versa so, only two answers!!! \[(a,b,c) = (0,\sqrt{2/3},-1/\sqrt{3})\qquad{\rm and}\\ (a,b,c) = (0,-\sqrt{2/3},1/\sqrt{3}) \]

OpenStudy (anonymous):

why, is it not sum of them =1

OpenStudy (anonymous):

not just sum, sum of squares!

OpenStudy (anonymous):

sum of square, so no matter how the sign is , they satisfy

OpenStudy (anonymous):

exactly. but that is not enough!

OpenStudy (anonymous):

check these results: (1) http://wolfr.am/12a2cVi (2) http://wolfr.am/XDGI4v

OpenStudy (anonymous):

ok, let it there, I'll be back to check, is it done?

OpenStudy (anonymous):

these are only two solutions. at the wolfram site, try to set them both positve and both negative and see if the result is still oorthogonal!